OpenStudy (mrhoola):

Question concerning exponential forms :

OpenStudy (mrhoola):

\[e ^{j2.5\pi + 2} = je ^{2}\]

OpenStudy (mrhoola):

I read over the chapter and wasnt able to find any information why this statement is true

OpenStudy (mrhoola):

I dont understand why it can be re written as je^2

OpenStudy (nubeer):

is it integration?

OpenStudy (mrhoola):

well , it is . here is the original problem .. one minute please

OpenStudy (mrhoola):

impulse function problem

hartnn (hartnn):

\(\Large e^{j2.5\pi }\times e^2 = (\sin 2.5\pi + j\cos 2.5\pi )e^2 = je^2\) doubts?

OpenStudy (anonymous):

typo... \[\cos\theta + i \sin \theta = e^{i\theta}\]

OpenStudy (unklerhaukus):

sin(2.5 π) = sin (0.5 π) cos(5/2 π) = cos(π/2)

OpenStudy (unklerhaukus):

\[e^{ j n π } = e^{ j (n±2) π }\]

hartnn (hartnn):

thanks for pointing it out! @Redcan \(\Large e^{j2.5\pi }\times e^2 = (\cos 2.5\pi + j\sin 2.5\pi )e^2 \\ \Large = (\cos \pi/2 + j\sin \pi/2 )e^2= je^2\)

OpenStudy (mrhoola):

@hartnn Wonderful ! Makes sense now.

OpenStudy (mrhoola):

Thanks everyone !